# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

### v1

@@ Acknowledgments
The data sets produced in this work can be obtained from the corresponding author upon an appropriate request.
@@
The proof utilizes induction. Initially, it is necessary to establish that the base case is valid at t=0. This statement is evidently valid given our initial assumption, particularly in the context of zero initialization. At a specific time-step t, we assume the validity of the statement, namely, $\sum_{z=0}^{Z}\theta_{t}^{z}=\theta_{t}^{\gamma}$. We will now demonstrate that it also holds at the subsequent time step, t+1.
$\sum_{z=0}^{Z}\theta_{t+1}^{z}$ $=\sum_{z=0}^{Z}\bigg{(}\theta_{t}^{z}+\alpha_{z}\bigg{(}G_{t}^{Z,\lambda_{z}}-\hat{W}_{z}(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t})\bigg{)}$
@@
$=\theta_{t}^{\gamma}+\sum_{z=0}^{Z}\alpha_{z}\bigg{(}\sum_{k=t}^{\infty}(\lambda_{z}\gamma_{z})^{k-t}\delta_{k}^{z}\bigg{)}\phi(s_{t},a_{t})\text{ , Assumption based on induction.}$.
$=\theta_{t}^{\gamma}+\alpha\sum_{k=t}^{\infty}(\lambda\gamma)^{k-t}\underbrace{\bigg{(}\sum_{z=0}^{Z}\delta_{k}^{z}\bigg{)}}_{*}\phi(s_{t},a_{t})\text{ , Because of $\alpha_{z}=\alpha$, $\lambda_{z}\gamma_{z}=\lambda_{\gamma}$, $\forall z$}$
To show that $\sum_{z=0}^{Z}\theta_{t+1}^{z}=\theta_{t+1}^{\gamma}$, we must establish that the term (*) = $\sum_{z=0}^{Z}\delta_{k}^{z}$ is equivalent to the conventional TD error $\delta_{k}^{\gamma}$.
$\sum_{z=0}^{Z}\delta_{k}^{z}$ $=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma_{0}}(s_{k},a_{k})+\sum_{z=1}^{Z}\bigg{(}(\gamma_{z}-\gamma_{z-1})\sum_{u=0}^{z-1}\langle\theta_{t}^{u},\phi(s_{k+1},a_{k+1})\rangle+\gamma_{z}\langle\theta_{t}^{z},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{t}^{z},\phi(s_{k},a_{k})\rangle\bigg{)}$
$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{z=1}^{Z}\bigg{(}\gamma_{z}\sum_{u=0}^{z}\theta_{t}^{u}-\gamma_{z-1}\sum_{u=0}^{z-1}\theta_{t}^{u}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{z=0}^{Z}\theta_{t}^{z},\phi(s_{k},a_{k})\bigg{\rangle}$
$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\gamma_{z}\bigg{\langle}\sum_{z=0}^{Z}\theta_{t}^{z},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\gamma_{0}\bigg{\langle}\theta_{t}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\gamma}(s_{k},a_{k})$
$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\gamma\hat{Q}_{\gamma}(s_{k+1},a_{k+1})-\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma}(s_{k},a_{k})$
$=r_{k}+\gamma\hat{Q}_{\gamma}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma}(s_{k},a_{k})$
$=\delta_{k}^{\gamma}$ (29)
@@
$\mathrm{T}=r+\gamma P$
The discount factor, the expected reward function, and the transition probability operator that is induced by the policy $\pi$ are denoted by the symbols $\gamma$, r, and P, respectively. A geometric weighted sum of $\mathrm{T}$ is the definition of the TD($\lambda$) operator, which can be expressed as follows:
@@
It may be inferred that $\lambda\in[0,1]$ is necessary for the aforementioned sum to be finite. A corresponding definition of $\mathrm{T}_{\lambda}$ is as follows for every function W:
$\mathrm{T}_{\lambda}W=W+(I-\lambda\gamma P)^{-1}(\mathrm{T}W-W)$ (32)
The separated action-value function is denoted by W in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation. In the event that $10\leq\lambda\gamma<$ 1, the formula is considered to be well-defined. This is because the spectral norm of the operator $\lambda\gamma P$ is less than 1, and thus, the Eq. $I-\lambda\gamma P$ is invertible. On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs. 31 and 32 is lost. This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq. 32. Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses. To begin, the Eq. designated as 32 can be rewritten as follows:
$\mathrm{T}_{\lambda}$ $=(I-\lambda\gamma P)^{-1}(\mathrm{T}W-\lambda\gamma PW)$ (33)
The contraction property is demonstrated by examining two action-value functions, $W_{1}$ and $W_{2}$. The distinction between the updates of $\mathrm{T}_{\lambda}W_{1}$ and $\mathrm{T}_{\lambda}W_{2}$ is:
$\mathrm{T}_{\lambda}W_{1}-\mathrm{T}_{\lambda}W_{2}$ $=(I-\lambda\gamma P)^{-1}\bigg{(}\mathrm{T}W_{1}-\mathrm{T}W_{2}-\lambda\gamma P(W_{1}-W_{2})\bigg{)}$
$=(I-\lambda\gamma P)^{-1}\bigg{(}\gamma P(W_{1}-W_{2})-\lambda\gamma P(W_{1}-W_{2})\bigg{)}$
$=(I-\lambda\gamma P)^{-1}\bigg{(}\gamma(1-\lambda)P(W_{1}-W_{2})\bigg{)}$
@@
$\|\mathrm{T}_{\lambda}W_{1}-\mathrm{T}_{\lambda}W_{2}\|\leq\frac{\gamma|1-\lambda|}{1-\lambda\gamma}\|W_{1}-W_{2}\|\textit{ where P = 1}$
@@
$\frac{\gamma(\lambda-1)}{1-\lambda\gamma}<1\Rightarrow\gamma\lambda-\gamma<1-\lambda\gamma$
$\Rightarrow 2\lambda\gamma<1+\gamma\Rightarrow\lambda<\frac{1+\gamma}{2\gamma}$ (34)
Consequently, $\mathrm{T}_{\lambda}$ constitutes a contraction if $0\leq\lambda<\frac{1+\gamma}{2\gamma}$. This directly indicates that for $\gamma<1$, $\gamma\lambda<1$.
@@
$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\gamma_{0}+\gamma r_{1}+...+\gamma^{k-1}r_{k-1}+\gamma^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached. The phased TD(k) update sums the terms $\gamma^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\gamma^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (35)
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (37)
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
@@
$\hat{Q}_{t+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}r_{0}+\gamma r_{1}+...+\gamma^{k-1}r_{k-1}+\gamma^{k}Q_{t}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
$=\sum_{i=0}^{k-1}\gamma^{i}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(j)}-\mathbb{E}[r_{i}]\bigg{)}+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (38)
@@
$\hat{Q}_{t+1}(s,a)\mathit{-}Q(s,a)\leq\sum_{i=0}^{k-1}\gamma^{l}\epsilon+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$
@@
$\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}$
$\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (39)
@@
$\gamma^{k}\Delta^{\hat{Q}_{\gamma}}_{t-1}$
and then the second term is bounded by $\Delta^{\hat{Q}_{\gamma}}_{t-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
$\Delta^{\hat{Q}_{\gamma}}_{t}\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\gamma^{k}\Delta^{\hat{Q}_{\gamma}}_{t-1}$ (40)
@@
Phased $TD(\Delta)$ update rules for $z\geq 1$:
$\hat{W}_{z,t}(s,a)=\frac{1}{n}\sum_{j-1}^{n}\bigg{(}\sum_{i=1}^{k_{z}-1}(\gamma_{z}^{i}-\gamma_{z-1}^{i})r_{i}^{(j)}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})\hat{Q}_{\gamma_{z-1}}(s_{k}^{(j)},a_{k}^{(j)})+\gamma_{z}^{k_{z}}\hat{W}_{z}(s_{k}^{(j)},a_{k}^{(j)})\bigg{)}$ (41)
We know that according to the multi-step update rule 14 for $z\geq 1$:
$W_{z}(s_{t},a_{t})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}r_{i}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}Q_{\gamma_{z-1}}(s_{k},a_{k})+\gamma_{z}^{k_{z}}W_{z}(s_{k},a_{k})\bigg{]}$ (42)
Then, subtracting the two expressions gives for $z\geq 1$:
$\hat{W}_{z,t}(s,a)-W_{z}(s_{t},a_{t})=\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\bigg{(}\sum_{u=0}^{z-1}\hat{W}_{u}(s_{k}^{(j)},a_{k}^{(j)})-\mathbb{E}[W_{z}(s_{k},a_{k})]\bigg{)}$
$+\gamma_{z}^{k_{z}}\bigg{(}W_{z}(s_{k}^{(j)},a_{k}^{(j)})-\mathbb{E}[W_{z}(s_{k},a_{k})]\bigg{)}$ (43)
Assume that $k_{0}\leq k_{1}\leq...k_{z}=k$, the W estimates share at most $k_{z}=k$ reward terms $\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. Hence, with probability $1-\delta$, $\forall z\geq 1$, we have:
$\Delta_{t}^{z}\leq\epsilon\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\epsilon\bigg{(}\frac{1-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}-\frac{1-\gamma_{z-1}^{k_{z}}}{1-\gamma_{z-1}}\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$ (44)
and $\Delta_{t}^{0}\leq\epsilon\frac{1-\gamma_{0}^{k_{0}}}{1-\gamma_{0}}+\gamma_{0}^{k_{0}}\Delta_{t-1}^{0}$
@@
$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\epsilon\frac{1-\gamma_{0}^{k_{0}}}{1-\gamma_{0}}+\epsilon\sum_{z=1}^{Z}\bigg{(}\frac{1-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}-\frac{1-\gamma_{z-1}^{k_{z}}}{1-\gamma_{z-1}}\bigg{)}+\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\underbrace{\epsilon\frac{1-\gamma_{Z}^{k_{Z}}}{1-\gamma_{Z}}+\epsilon\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}}_{\text{(*)variance term}}+\underbrace{\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}}_{\text{(**)bias term}}$ (45)

### v2

@@ Acknowledgments
The proof utilizes induction. Initially, it is necessary to establish that the base case is valid at n=0. This statement is evidently valid given our initial assumption, particularly in the context of zero initialization. At a specific time-step n, we assume the validity of the statement, namely, $\sum_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$. We will now demonstrate that it also holds at the subsequent time step, n+1.
$\sum_{m=0}^{M}\theta_{n+1}^{m}$ $=\sum_{m=0}^{M}\bigg{(}\theta_{n}^{m}+\alpha_{m}\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n})\bigg{)}$
@@
$=\theta_{n}^{\eta}+\sum_{m=0}^{M}\alpha_{m}\bigg{(}\sum_{k=n}^{\infty}(\lambda_{m}\eta_{m})^{k-n}\delta_{k}^{m}\bigg{)}\phi(s_{n},a_{n})\text{ , Assumption based on induction.}$.
$=\theta_{n}^{\eta}+\alpha\sum_{k=n}^{\infty}(\lambda\eta)^{k-n}\underbrace{\bigg{(}\sum_{m=0}^{M}\delta_{k}^{m}\bigg{)}}_{*}\phi(s_{n},a_{n})\text{ , Because of $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$,$\forall m$}$
To show that $\sum_{m=0}^{M}\theta_{n+1}^{m}=\theta_{n+1}^{\eta}$, we must establish that the term (*) = $\sum_{m=0}^{M}\delta_{k}^{m}$ is equivalent to the conventional TD error $\delta_{k}^{\eta}$.
$\sum_{m=0}^{M}\delta_{k}^{m}$ $=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta_{0}}(s_{k},a_{k})+\sum_{m=1}^{M}\bigg{(}(\eta_{m}-\eta_{m-1})\sum_{q=0}^{m-1}\langle\theta_{n}^{q},\phi(s_{k+1},a_{k+1})\rangle+\eta_{m}\langle\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{n}^{m},\phi(s_{k},a_{k})\rangle\bigg{)}$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{m=1}^{M}\bigg{(}\eta_{m}\sum_{q=0}^{m}\theta_{n}^{q}-\eta_{m-1}\sum_{q=0}^{m-1}\theta_{n}^{q}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k},a_{k})\bigg{\rangle}$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta_{m}\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\eta_{0}\bigg{\langle}\theta_{n}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\eta}(s_{k},a_{k})$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
$=r_{k}+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
$=\delta_{k}^{\eta}$ (32)
@@
$\mathrm{T}=r+\eta P$
The discount factor, the expected reward function, and the transition probability operator that is induced by the policy $\pi$ are denoted by the symbols $\eta$, r, and P, respectively. A geometric weighted sum of $\mathrm{T}$ is the definition of the TD($\lambda$) operator, which can be expressed as follows:
@@
It may be inferred that $\lambda\in[0,1]$ is necessary for the aforementioned sum to be finite. A corresponding definition of $\mathrm{T}_{\lambda}$ is as follows for every function D:
$\mathrm{T}_{\lambda}D=D+(I-\lambda\eta P)^{-1}(\mathrm{T}D-D)$ (35)
The separated action-value function is denoted by D in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation. In the event that $10\leq\lambda\eta<$ 1, the formula is considered to be well-defined. This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, the Eq. $I-\lambda\eta P$ is invertible. On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs. 34 and 35 is lost. This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq. 35. Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses. To begin, the Eq. designated as 35 can be rewritten as follows:
$\mathrm{T}_{\lambda}$ $=(I-\lambda\eta P)^{-1}(\mathrm{T}D-\lambda\eta PD)$ (36)
The contraction property is demonstrated by examining two action-value functions, $D_{1}$ and $D_{2}$. The distinction between the updates of $\mathrm{T}_{\lambda}D_{1}$ and $\mathrm{T}_{\lambda}D_{2}$ is:
$\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}$ $=(I-\lambda\eta P)^{-1}\bigg{(}\mathrm{T}D_{1}-\mathrm{T}D_{2}-\lambda\eta P(D_{1}-D_{2})\bigg{)}$
$=(I-\lambda\eta P)^{-1}\bigg{(}\eta P(D_{1}-D_{2})-\lambda\eta P(D_{1}-D_{2})\bigg{)}$
$=(I-\lambda\eta P)^{-1}\bigg{(}\eta(1-\lambda)P(D_{1}-D_{2})\bigg{)}$
@@
$\|\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}\|\leq\frac{\eta|1-\lambda|}{1-\lambda\eta}\|D_{1}-D_{2}\|\textit{ where P = 1}$
@@
$\frac{\eta(\lambda-1)}{1-\lambda\eta}<1\Rightarrow\eta\lambda-\eta<1-\lambda\eta$
$\Rightarrow 2\lambda\eta<1+\eta\Rightarrow\lambda<\frac{1+\eta}{2\eta}$ (37)
Consequently, $\mathrm{T}_{\lambda}$ constitutes a contraction if $0\leq\lambda<\frac{1+\eta}{2\eta}$. This directly indicates that for $\eta<1$, $\eta\lambda<1$.
@@
$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\eta_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached. The phased TD(k) update sums the terms $\eta^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\eta^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (38)
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (40)
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
@@
$\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}r_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q_{n}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
$=\sum_{i=0}^{k-1}\eta^{i}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (41)
@@
$\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)\leq\sum_{i=0}^{k-1}\eta^{l}\epsilon+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$
@@
$\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}$
$\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (42)
@@
$\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$
and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
$\Delta^{\hat{Q}_{\eta}}_{n}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$ (43)
@@
Phased $TD(\Delta)$ update rules for $m\geq 1$:
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\bigg{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (44)
We know that according to the multi-step update rule 17 for $m\geq 1$:
$D_{m}(s_{n},a_{n})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}r_{i}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}Q_{\eta_{m-1}}(s_{k},a_{k})+\eta_{m}^{k_{m}}D_{m}(s_{k},a_{k})\bigg{]}$ (45)
Then, subtracting the two expressions gives for $m\geq 1$:
$\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$
$+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$ (46)
Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. Hence, with probability $1-\delta$, $\forall m\geq 1$, we have:
$\Delta_{n}^{m}\leq\epsilon\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\epsilon\bigg{(}\frac{1-\eta_{m}^{k_{m}}}{1-\eta_{m}}-\frac{1-\eta_{m-1}^{k_{m}}}{1-\eta_{m-1}}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$ (47)
and $\Delta_{n}^{0}\leq\epsilon\frac{1-\eta_{0}^{k_{0}}}{1-\eta_{0}}+\eta_{0}^{k_{0}}\Delta_{n-1}^{0}$
@@
$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\frac{1-\eta_{0}^{k_{0}}}{1-\eta_{0}}+\epsilon\sum_{m=1}^{M}\bigg{(}\frac{1-\eta_{m}^{k_{m}}}{1-\eta_{m}}-\frac{1-\eta_{m-1}^{k_{m}}}{1-\eta_{m-1}}\bigg{)}+\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\underbrace{\epsilon\frac{1-\eta_{M}^{k_{M}}}{1-\eta_{M}}+\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta_{m}^{k_{m}}}{1-\eta_{m}}}_{\text{(*)variance term}}+\underbrace{\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}}_{\text{(**)bias term}}$ (48)